OpenStudy (anonymous):

A company invests $15,000.00 in an account that compounds interest annually. After two years, the account is worth $16,099.44. Use the function in which r is the annual interest rate, P is the principal, and A is the amount of money after t years. What is the interest rate of the account? A = P(1 + r)t 1.04% 3.6% 5.4% 7.3%

OpenStudy (anonymous):

The definition of an extraneous solution is one that emerges as a result of solving the problem, even though it isn't valid. None of these solutions qualify. Sorry, but let me try to help you with #2. Here is the present value equation that we'll use to solve for the interest rate: PV=FV/(1+r)^t. PV is the present value of the money, FV is the future value, r is the rate of interest, and t is the time period. Thus, 16,099.44/(1+r)^2=15,000; 1.073=(1+r)^2; 1.036=1+r; r=.036=3.6%. Hope this helps.

OpenStudy (anonymous):

so did you ever get the answer?

OpenStudy (anonymous):

It's says right there... 3.6% @ShadowFang

OpenStudy (anonymous):

so its 4 right?

OpenStudy (anonymous):

4%*

OpenStudy (anonymous):

ANSWER: 3.6%